a ring is made of metal gold and silver the weights of the metals are in the ratio 8:9. if the weight if the ring is 1/8 of a pund then what are the weights of the silver and gold?
Answer by mananth(16946)

ratio of gold to silver = 8:9
There are totally 17 parts ofgold and silver
8/17 is silver and 9/17 part is gold
total weight of ring = 1/8 of pound
gold will be 8/17 * 1/8
= 1/17 pound
and silver will be 9/17 *1/8
=9/ 136 of pound
Ananth